OpenStudy (anonymous):

The sum of the integers p and q is 495. If p is divided by 10, the result is equal to q. What is the value of p ? a.40 b.45 c.245 d.250 e.450

OpenStudy (anonymous):

p + q = 495 p = 10q substitution. 10q + q = 495 q = 45 p = 10q = 10(45) = 450

OpenStudy (anonymous):

where did the 45 come from

OpenStudy (anonymous):

11q = 495 q = 495/11

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

glad i could help :)